Suppose f(x,y)=(x−y)(16−xy). Find the saddle points of f.
1
Expert's answer
2020-04-06T16:07:15-0400
"f(x, y)=(x-y)(16-xy)=16x-x^2y-16y+xy^2"

Find the critical point(s)


"f_x=16-2xy+y^2, f_y=-16+2xy-x^2"

"\begin{cases}
 f_x=0 \\
 f_y=0
\end{cases}=> \begin{cases}
 16-2xy+y^2=0 \\
 -16+2xy-x^2=0
\end{cases}=>""=> \begin{cases}
 (y-x)(y+x)=0 \\
 -16+2xy-x^2=0
\end{cases}"

Critical points: "(-4,-4),\ (-4,4),\ (4,-4),\ (4,4)"


Second Derivatives Test


"f_{xx}=-2y, f_{xy}=f_{yx}=-2x+2y, f_{yy}=2x"


"D=\begin{vmatrix}
 f_{xx} & f_{xy} \\
 f_{yx} & f_{yy}
\end{vmatrix}=\begin{vmatrix}
 -2y & -2x+2y \\
 -2x+2y & 2x
\end{vmatrix}="

"=-4xy-4y^2 +8xy-4x^2=-4(x^2-xy+y^2)"


"x^2-xy+y^2>0, x\in\R, y\in \R=>"

"=>D(x,y)<0, x\in\R, t\in \R"

All critical points are saddle points.

"(-4,-4):"


"f(-4,-4)=(-4-(-4))(16-(-4)(-4))=0"

"(-4,4):"


"f(-4,4)=(-4-4)(16-(-4)(4))=-256"

"(4,-4):"


"f(4,-4)=(4-(-4))(16-4(-4))=256"

"(4,4):"

"f(4,4)=(4-4)(16-4(4))=0"

Saddle points: "(-4,-4,0),\ (-4,4,-256),\ (4,-4,256),\ (4,4,0)"